OpenStudy (anonymous):

A right circular cylinder has a total surface area of 112pie. If the height of the cylinder is 10, find the diameter of the base.

OpenStudy (anonymous):

Radius = 4 units and Diameter = 8 units

OpenStudy (anonymous):

How is this problem worked out though?

OpenStudy (anonymous):

formula is Surface Area of RCC = 2*Pie* r *(h+r) 112* Pie = 2* Pie * (10r + r^2) 56 = 10r + r^2 on solving u will get r= 4 and r = -10 but r can't be negative so r = 4 d = 8